OpenStudy (anonymous):

Use complete sentences to describe the steps taken to simplify this problem. Make sure you include information about the new common denominator, the final simplified expression, and any restrictions. 3 over x plus the fraction 3 over x plus 1 plus the fraction 3 over x plus 2.

OpenStudy (anonymous):

\[\frac{ 3 }{ x } + \frac{ 3 }{ x+1 } +\frac{ 3 }{ x+2 }\]